OpenStudy (anonymous):

hi-- G(w)= 3+2/w-1. is G(w) = -1 in range? I think it should be, but to find the domain I plug in -1 and am getting w=-1. The book gets 1/2.

OpenStudy (turingtest):

I assume you mean to have parentheses like G(w)=3+2/(w-1) correct?

OpenStudy (turingtest):

\[G(w)=3+2/(w-1)\]?????????????

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

any thoughts?

OpenStudy (turingtest):

ok so if -1 is in the range we must find some w such that G(w)=-1=3+2/(w-1) -4=2/(w-1) w-1=-1/2 w=1/2 so there is the answer. You say "to find the domain I plug in -1 and am getting w=-1" which is not right, you seem confused. The Domain is the set of all w for which G(w) is defined, so the Domain is \[w \neq 1\]because when w=1 we get\[G(w)=3+2/0=undefined\]The Range are all possible values of the function G(w).